Solve for R
0=-16250000+250000000(.09+R)/R

Answer by Alan3354(69443)

Solve for R
0=-16250000+250000000(.09+R)/R
---------
Multiply by R
-16250000+250000000(.09+R) = 0
Add 16250000
250000000(.09+R) = 16250000
Divide by 10000
25000(.09+R) = -1625
divide by 25000
0.09+R = -0.065
Subtract 0.09
R = -0.155